Plspls:

A car running with uniform speed, covers a distance of 98km in 3 hrs. How much distance will the car cover in 5 hrs, running with the same speed?

MysticalDimpless:

In 3 hours, car covers 98 km → In 1 hour, car covers 96/3 km = 32 km And, In 5 hours, car covers 32km x 5 = 160km

Plspls:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @MysticalDimpless In 3 hours, car covers 98 km → In 1 hour, car covers 96/3 km = 32 km And, In 5 hours, car covers 32km x 5 = 160km \(\color{#0cbb34}{\text{End of Quote}}\) Ty

XioGonz:

3 hours = 96 km 1 hour = (96/3) Km/h 1 hour = 32 km/h 5 hours = (32 x 5) Km 5 hours = 160 Km

bryna321:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @XioGonz 3 hours = 96 km 1 hour = (96/3) Km/h 1 hour = 32 km/h 5 hours = (32 x 5) Km 5 hours = 160 Km \(\color{#0cbb34}{\text{End of Quote}}\) @MysticalDimpless already answered lol

XioGonz:

Yes. My work is origin though. Not taken off of brainly as that answer is.

XioGonz:

Original*

MysticalDimpless:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @XioGonz Yes. My work is origin though. Not taken off of brainly as that answer is. \(\color{#0cbb34}{\text{End of Quote}}\) My answer is not from brainly

XioGonz:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @MysticalDimpless \(\color{#0cbb34}{\text{Originally Posted by}}\) @XioGonz Yes. My work is origin though. Not taken off of brainly as that answer is. \(\color{#0cbb34}{\text{End of Quote}}\) My answer is not from brainly \(\color{#0cbb34}{\text{End of Quote}}\) The answer you've posted is very similar to the format.

XioGonz:

I apologize if I am wrong though.

MysticalDimpless:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @XioGonz \(\color{#0cbb34}{\text{Originally Posted by}}\) @MysticalDimpless \(\color{#0cbb34}{\text{Originally Posted by}}\) @XioGonz Yes. My work is origin though. Not taken off of brainly as that answer is. \(\color{#0cbb34}{\text{End of Quote}}\) My answer is not from brainly \(\color{#0cbb34}{\text{End of Quote}}\) The answer you've posted is very similar to the format. \(\color{#0cbb34}{\text{End of Quote}}\) Ik but it is not copied as you think

XioGonz:

I'll try to take your word for it. I'm just letting you know now. Here we don't copy solutions from other websites. Our solutions have to be completely original and our own. If not, then it is considered plagiarism, something of which isn't tolerable in question cove.

MysticalDimpless:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @XioGonz I'll try to take your word for it. I'm just letting you know now. Here we don't copy solutions from other websites. Our solutions have to be completely original and our own. If not, then it is considered plagiarism, something of which isn't tolerable in question cove. \(\color{#0cbb34}{\text{End of Quote}}\) Wdym? I didnt copied anything it is original then how is it plagiarism?

XioGonz:

As I said earlier, I'll try to take your word for it.

MysticalDimpless:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @XioGonz As I said earlier, I'll try to take your word for it. \(\color{#0cbb34}{\text{End of Quote}}\) Okay

XioGonz:

@jhonyy9

jhonyy9:

A car running with uniform speed, covers a distance of 98km in 3 hrs. How much distance will the car cover in 5 hrs, running with the same speed? so first of all i dont know from where you get this 96/3 = 32 just bc. there is given 98 km in 3 hours not 96 km

jhonyy9:

so using these given details and the rule of 3 simple we can write 98 km covers in 3 hrs x km covers in 5 hrs - cross multiplie and calcule the x value x = 98*5/3 = ? km hope these help you understandably

jhonyy9:

@XioGonz

XioGonz:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @jhonyy9 so using these given details and the rule of 3 simple we can write 98 km covers in 3 hrs x km covers in 5 hrs - cross multiplie and calcule the x value x = 98*5/3 = ? km hope these help you understandably \(\color{#0cbb34}{\text{End of Quote}}\) Thank You jhony I didn't realize my mastake. Thank you for the correction.

jhonyy9:

np anytime